Please Help!
The question is:
The Sum of two numbers is 25. Tweleve less than four times one of the numbers is 16 more than twice the other number. Find both numbers.
I tried figuring it out, but I got that 12&39=25, which is not possible. Please help!

Answer by checkley71(8403)

X+Y=25 OR X=25-Y
4X-12=2Y+16 NOW SUBSTITUTE (25-Y) FOR X IN THE SECOND EQUATION & SOLVE FOR Y:
4(25-Y)-12=2Y+16
100-4Y-12=2Y+16
100-16-12=2Y+4Y
6Y=72
Y=72/6
Y=12 FOR ONE NUMBER.
X+12=25
X=25-12
X=13 FOR THE SECOND NUMBER
PROOF
4*13-12=2*12+16
52-12=24+16
40=40